OpenStudy (anonymous):

PLEASE HELP Two angles are complementary. One angle is 4 degrees less than three times the other angle. Find the measure of each angle

OpenStudy (mertsj):

Live preview of your reply: Let x = the angle 3x-4 = the complement Add the two angles and set the sum equal to 90 Solve the resulting equation.

OpenStudy (radar):

@mimi0921 Just add those two angles provided by Mersj.

OpenStudy (radar):

3x - 4 x add ------------

OpenStudy (radar):

That resulted in ?? What did you get??

OpenStudy (anonymous):

4X-4? I have no idea. I'm not understanding it at all.

OpenStudy (radar):

Yes that is the sum of the two angles. Now focus here. The two angles are also complimentary. So what would the sum be equal to??

OpenStudy (anonymous):

90 ?

OpenStudy (radar):

Absolutely, can you now solve for x and 3x-4 ?? Use algebra. 4x - 4 = 90

OpenStudy (anonymous):

x= 23.5

OpenStudy (radar):

Yes, and the other one?

OpenStudy (ahsome):

Ok, lets do this slowly. When they say two angles are complimentary, they mean that when you add them up, they equal 90. Example: Angle A is 40 Degrees Angle B is 50 Degrees We can say Angle A and Angle B are complimentary, since they add up to 90 Now, in your question, there are two angles. Lets make the first angle \(x\) and the second one \(y\) They are also saying that \(y\) is equal to is four less than three times angle \(x\). So, we can say: \(x=x\) (Obviously) \(y=3x-4\) Now, since they are complimentary, they add up to 90: \(x+y=90\) Lets put our past \(y\) value: \(x+3x-4=90\) Now, solve for \(x\) \(x+3x=94\) \(4x=94\) \(x=\frac{94}{4}\) So, angle \(x\) is 23.5 degrees. What you can do now is put that value to: \[y=3x-4\] And find angle \(y\)

OpenStudy (anonymous):

x=31.3

OpenStudy (radar):

Substitute that value for x in 3x-4 or 3(23.5) - 4 or 66.5

OpenStudy (radar):

Or since you know the sum is 90 you could just subtract the 23.5 from 90. 66.5 degrees

OpenStudy (radar):

Ahsome has done an excellent series of steps that reveal the method also.

OpenStudy (anonymous):

so x=23.5 and y=66.5 ?

OpenStudy (anonymous):

I REALLY APPRECIATE BOTH OF YALLS HELP IN HELPING ME!!

OpenStudy (radar):

Yes, when you add them up they are indeed 90 degrees. No problem good luck with your studies.

OpenStudy (anonymous):

thank you !

OpenStudy (ahsome):

No problem @mimi0921 :D